Show that the plane polarized light and circularly polarized light are the special cases of elliptically polarized light
1
Expert's answer
2020-02-18T09:36:46-0500

In general, projections of light vectors of polarized light on the axes X and Y


"E_x=E_0\cos(\omega t-kz)"


"E_y=E_0\cos(\omega t-kz+\phi)"


they satisfy the equation


"\frac{E^2_x}{E^2_0}-2\frac{E_xE_y}{E^2_0}\cos\phi+\frac{E^2_y}{E^2_0}=\sin^2\phi"


This equation is an equation of an ellipse whose axes are oriented relative to the coordinate axes X and Y arbitrarily. The orientation of the ellipse and magnitude of its semiaxes depends only on the angle "\phi" (phase difference).


if "\phi=\pi"


"E_x+E_y=0"


we have the equation of the line. That is, light is linearly or plane polarized.


If "\phi=0"


"E^2_x+E^2_y=E^2_0"


we have the equation of the circle. That is, light is circularly polarized.